### v1

@@ Abstract
A high-resolution hindcast archive around the coast of Norway spanning the time period 1979-2022 is presented. The model is found to perform well with forcing from the NORA3 atmospheric hindcast. Observations from water level gauges along the coasts of the North Sea and the Norwegian Sea are used to assess the performance of the hindcast, which is generally found to be in good agreement. In addition, 100-year return level estimates are calculated from the hindcast.
@@ 1 Introduction
Storm surges are driven by the passing of synoptic weather systems. A simple rule of thumb states that a reduction of 1 hPa in surface air pressure leads to an increase of 1 cm in water level from the inverse barometric effect alone. In addition, frictional forces will lead to Ekman transport which can cause convergence and divergence of water masses along coastlines. More generally, such perturbations of the sea surface will tend to manifest themselves as Kelvin waves, travelling with the coast to their right in the northern hemisphere. As these waves have spatial and temporal scales comparable to the astronomical tides, they have historically been known as tidal surges, but their generating mechanism is of course very different. More importantly, storm surges are only as predictable as the synoptic weather phenomena that cause them. There exists a number of well-documented water level forecast systems in Europe and elsewhere. The simplest represent the ocean as a two-dimensional barotropic field with a free surface. This is also how we model the water level in this study. As Kelvin waves propagate as shallow-water waves whose phase and group speed are $\sqrt{gH}$, where $H$ is the local depth and $g\approx 9.81\,\mathrm{m\,s^{-2}}$ the earth’s gravitational acceleration, the amplitude and propagation speed of storm surges is sensitive to errors in both the gradient and the mean level of the bathymetric field.
Long hindcast integrations are necessary to assess water level extremes in coastal regions. To do this realistically requires high-quality and high-resolution atmospheric forcing and sufficiently high spatial resolution for the storm surge ocean model. Storm surge hindcast archives serve mainly three practical applications. The first is to properly account for the overall statistical distribution of the water level, and the other is to assess the extreme value distribution, i.e., the return values for extreme high and low water level. In addition to these statistical measures, the scarcity of observational records means that accurate reconstructions of past events are of societal value in themselves. The objective of this article is to present such a historical hindcast and to assess its validity for extreme value estimation of water level in the North Sea, the Norwegian Sea and the Barents Sea.
@@ 2 Observational dataset
Water level gauges from the coast of Norway (stations with NO prefix, also used in the study by Tedesco et al. 2023), Norwegian oil platforms (WL), Denmark (DK), Sweden (SW), Germany (DE), the Netherlands (NL), and the United Kingdom (UK) make up our observational data. Unfortunately, the license agreements do not allow open redistribution of the data we gathered. However, we share the code used for parsing and assembling the data from these sources into a single dataset (see Appendix A). We also summarize the source of the data from each country in Appendix B.
The data provided come from a variety of water level gauges. The location of the gauges and the associated measurement stations are shown in Figure 1. The time spans of the time series vary from station to station. Some stations have continuous recordings going back to 1915 (for some of the UK stations), though our dataset is built starting from 1979 onwards. As a consequence of the large time span covered, the kind of gauge, temporal resolution, and water level resolution of the measurements is not homogeneous in time even sometimes for a single individual station. Moreover, some data sources provide a quality flag as part of the data downloaded (depending on the data source and station, this may be either an automatic quality flag, or a human determined expert quality flag), while others only provide raw data without quality flag. We have, therefore, performed a detailed quality control cycle on all the data, which includes quality checking of individual measurements through both flag checking and statistical analysis, interpolation on a common time base, detrending, visual inspection, automated outlier detection, and a series of quality checks on running windows of observational data. The technical details of how this was performed are presented in Appendix B.
A tidal analysis was performed with the CNES pangeo-pytide python package  [^1] https://github.com/CNES/pangeo-pytide in order to estimate the tidal signal at each station and to detide the observed water level. For this, the time series obtained for each station is used to obtain the modal coefficients for the default 67 tidal modes used by pytide 222the corresponding modes are: O1, P1, K1, 2N2, Mu2, N2, Nu2, M2, L2, T2, S2, K2, M4, S1, Q1, Mm, Mf, Mtm, Msqm, Eps2, Lambda2, Eta2, 2Q1, Sigma1, Rho1, M11, M12, Chi1, Pi1, Phi1, Theta1, J1, OO1, M3, M6, MN4, MS4, N4, R2, R4, S4, MNS2, M13, MK4, SN4, SK4, 2MN6, 2MS6, 2MK6, MSN6, 2SM6, MSK6, MP1, 2SM2, Psi1, 2MS2, MKS2, 2MN2, MSN2, MO3, 2MK3, MK3, S6, M8, MSf, Ssa, Sa, and to perform tidal prediction on the same time base as the observations.
@@ 3.1 The ocean model
The Regional Ocean Model System (ROMS, see Shchepetkin and McWilliams 2005) is used for operational storm surge forecasting for Norwegian waters, as described by Kristensen et al. (2022). The model is set up to run in two-dimensional barotropic mode and the model domain is depicted in Figure 2. The governing model equations are thus the shallow water equations as described in Haidvogel et al. (2008). Let the total height be $h=H+\zeta$, where the $H$ is the equilibrium depth and $\zeta$ is the sea surface deviation from the mean. By averaging the velocity over the total height $h$ we get (Kristensen et al., 2022; Tedesco et al., 2023),
@@ 4.1 Model performance evaluated against observations
The hindcast dataset has been extensively evaluated and validated against the observational dataset obtained in Section 2. In all the following analyses, the mean sea level for the entire hindcast period for each grid point has been subtracted from the modeled water level in order to establish mean sea level as the reference level.
@@
Figure 3: Left panel: Comparison of standard deviation between hindcast and observed time series for each station. In addition, the stations that were disregarded in the analysis are marked by triangles. Right panel: Difference between model and observation for all stations for the entire hindcast period. Note that the temporal coverage of the observing stations varies significantly. Stations are grouped by country.
@@
Figure 3a shows a comparison of the standard deviation for the time series for both the hindcast data and the observational data for all stations. This can be viewed as a measure of storm surge activity, and clearly shows that the hindcast data exhibit similar properties with regards to activity as the observations, and hence confirms the general quality of the hindcast. In Figure 3b we present the difference between the hindcast and observational data for each station.
MAE [cm] RMSE [cm]
All data 9.72 12.39
Masked less than +/- 10cm 10.07 12.88
Masked less than +/- 30cm 11.82 15.19
@@
Table 1: General values for Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) averaged over all stations over the entire hindcast period. The rows with "masked" values indicate MAE and RMSE values when masking the data for the time periods when the observed absolute values of storm surge is less than $10$ and $30$ cm. The masking thresholds removes $43\%$ and $85\%$ of the data, respectively.
@@
Average values for Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) for all stations combined, averaged over the entire hindcast period, are presented in Table 1. When including all the available quality controlled observations in the comparison with the model data we get a MAE of $9.72$ cm and RMSE of $12.39$ cm. On average, the storm surge contribution to total water level is relatively small, as can be seen in the histograms in Figure 4, showing the distribution of modeled and observed water level, and in Figure 5, which displays the density scatterplot of observations vs modeled values. To investigate the events when there is actually a significant contribution from storm surge to the total water level, we have also added the values of MAE and RMSE where we have masked the data when the observed absolute value of the amplitude of storm surge is less than $10$ cm and $30$ cm in Table 1. These thresholds remove $43\%$ and $85\%$ of the data, respectively. As expected, and as explained by Kristensen et al. (2022), when the amplitudes of storm surge increase, the MAE and RMSE are expected to increase.
@@
The distributions of modeled and observed water level in Figure 4 indicate that the hindcast tends to slightly underestimate the storm surge in the range $-50$ to $+20$ cm when considering all stations combined (Figure 4a) and the range ±$\pm 25$ cm for the Norwegian stations (Figure 4b). For the range extending further above $+25$ cm, the hindcast has a slight increase in count for values up to about $+200$ cm for all stations. When examining the details of the tail of the distribution, for both positive and negative water level in Figures 4c and 4d, we note that the hindcast generally does not contain the same extremal values as the observations. Also, we would like to point out the four outliers with values below $-400$ cm and above $+400$ cm in Figure 4c. These are believed to be unphysical outliers in the observational dataset that evaded our quality control.
@@
The heatmaps displaying the two-dimensional distribution of corresponding hindcast and observation values in Figure 5 further confirm that the hindcast is in good agreement with the observations. This is borne out by the quantitative statistics presented in Table 1.
@@
We note a small tendency for observations to be clustered around zero in the distribution for all stations (Figure 5a). The same is seen, although less pronounced, for the hindcast, with a slight increase in density along the $x$-axis. This could be caused by either unrealistic or erroneous observations still left in the dataset after quality control, or wrong characteristics for the hindcast for some stations/periods. The same is not seen in the distribution for the Norwegian stations (Figure 5b). Further investigation into the subject narrows the feature down to the stations located in the Netherlands (most pronounced), Germany and Denmark. Since the number of datapoints with this behaviour is relatively small (note the log scale for density), further efforts to identify and possibly remove these data points/stations has not been made. This might also be an indication that the NORA-Surge hindcast in some cases has problems capturing the exact behaviour of the storm surge for shallow stations located in the German Bight and surrounding area where tide-surge interaction might be of importance.
Our model setup is tuned to, and targeted at, providing good hindcasts along the Norwegian coast. Because of this, we expect some geographical differences in the quality of the hindcast. These geographical differences can also be attributed to the fact that storm surges behave differently in different geographical areas due to, e.g., different characteristics of bottom topography, coastline geometry and weather patterns. Some of these differences can be seen in Figure 3b where we display the difference between model and observations for all stations (grouped by country) for the entire time coverage of the hindcast. By comparing the standard deviation in Figure 3a with the differences in Figure 3b, there is a clear pattern where the most saturated red and blue colors (i.e., the largest differences) occur at the stations with the highest standard deviations. In Table 2 we provide the same type of comparison as in Table 1, but grouped by country. The smallest absolute errors are found at the stations in Sweden, Norway and the offshore oil platforms located in the North Sea, with values of MAE and RMSE, respectively, of less than $7$ cm and $9$ cm. This is less than the averages over all stations, and more in line with the values of MAE and RMSE of $3$ and $5$ cm for the errors reported at initial time ($+0$ h forecast lead time) by Kristensen et al. (2022) for the Norwegian storm surge forecast system evaluated at Norwegian stations. However, we would like to point out that the forecasts used in the evaluation by Kristensen et al. (2022) are post-processed and subject to correction by using the difference between model and observations, and hence exhibit lower error values than the free run in the NORA-Surge hindcast. The purpose of such corrections is to reduce low frequency error components in the reference level for the water level present in the operational forecast for a discrete set of stations on the Norwegian Coast. The correction does not affect the amplitude of storm surges, and does not correct the two-dimensional model field. Thus, such a correction is not applicable along the whole coastline in the present hindcast.
MAE RMSE
Denmark All data 10.70 13.67
Masked less than +/- 10cm 10.88 14.03
Masked less than +/- 30cm 12.21 15.96
Germany All data 13.88 17.80
Masked less than +/- 10cm 14.28 18.46
Masked less than +/- 30cm 16.17 21.06
Netherlands All data 11.65 14.99
Masked less than +/- 10cm 11.95 15.40
Masked less than +/- 30cm 13.81 17.79
Norway All data 6.94 8.56
Masked less than +/- 10cm 7.22 8.94
Masked less than +/- 30cm 8.49 10.40
Sweden All data 6.52 8.27
Masked less than +/- 10cm 6.82 8.66
Masked less than +/- 30cm 8.23 10.28
UK All data 8.45 10.71
Masked less than +/- 10cm 9.10 11.49
Masked less than +/- 30cm 11.20 13.99
Oil platforms All data 5.74 7.58
Masked less than +/- 10cm 6.42 8.71
Masked less than +/- 30cm 9.93 14.90
@@
We note two outliers in the Taylor diagram, the top left NO station with a correlation of approximately $0.7$ and the UK station with a correlation of $0.75$. The NO station is Ny-Ålesund (NO_NYA) located on the Arctic island of Spitsbergen, and the UK station is Newhaven (UK_Newhaven), located just west of Dover. For the NO_NYA station, we note that there seems to be a negative trend in the difference between hindcast and observation in Figure 3b. This could indicate an issue with the observations for this station, and we considered removing the station from the dataset, but chose to still include it due to its unique location far north in the Arctic. Other possible explanations for the reduced quality for this station could be effects of sea ice not present in the model, or the fact that, even with its position near the open boundary of the model domain, this station has the longest potential travel distance for storm surges in our model domain. Any errors due to imperfections in the model bottom bathymetry will accumulate as the typical storm surge Kelvin wave moves northward through the Norwegian Sea. The UK_Newhaven station is placed near the south-western boundary of the hindcast model domain, and the fact that it is located west of Dover means it will be subject to storm surges that enter the English Channel from the west. The lack of proper boundary values from an outer model, as mentioned in Section 3.1, could cause some of the storm surge signal that should have entered from the North Atlantic to be missing in the hindcast for this area of the English Channel.

### v2

@@ Abstract
Knowledge about statistics for water level variations along the coast due to storm surge is important for the utilization of the coastal zone. An open and freely available storm surge hindcast archive covering the coast of Norway and adjacent sea areas spanning the time period 1979-2022 is presented. The storm surge model is forced by wind stress and mean sea level pressure taken from the non-hydrostatic NORA3 atmospheric hindcast. A dataset consisting of observations of water level from more than 90 water level gauges along the coasts of the North Sea and the Norwegian Sea is compiled and quality controlled, and used to assess the performance of the hindcast. The observational dataset is distributed in both time and space, and when considering all the available quality controlled data, the comparison with modelled water levels yield a mean absolute error (MAE) of $9.7$ cm and a root mean square error (RMSE) of $12.4$ cm. Values for MAE and RMSE scaled by the standard deviation of the observed storm surge for each station are $0.42$ and $0.54$ standard deviations, repsectively. When considering the geographical differences in characteristics of storm surge for different countries/regions, the values of MAE and RMSE are in the range $5.7-13.9$ cm and $7.6-17.8$ cm respectively, and $0.33-0.46$ and $0.42-0.59$ standard deviations for the scaled values. The minimum and maximum values for water level in the hindcast are $-2.60$ m and $3.92$ m. In addition, 100-year return level estimates are calculated from the hindcast, with minimum and maximum values of, respectively, $-2.75$ m and $3.98$ m. All minimum and maximum values are found in the southern North Sea area.
@@ 1 Introduction
Storm surges are driven by the passing of synoptic weather systems. A simple rule of thumb states that a reduction of 1 hPa in surface air pressure leads to an increase of 1 cm in water level from the inverse barometric effect alone. In addition, frictional forces in combination with the Coriolis effect will lead to Ekman transport which can cause convergence and divergence of water masses along coastlines. More generally, such perturbations of the sea surface will tend to manifest themselves as Kelvin waves, travelling with the coast to their right in the northern hemisphere. As these waves have spatial and temporal scales comparable to the astronomical tides, they have historically been known as tidal surges, but their generating mechanism is of course very different. More importantly, storm surges are only as predictable as the synoptic weather phenomena that cause them. There exists a number of well-documented water level forecast systems in Europe and elsewhere (e.g. Glahn et al. 2009; Zijl et al. 2013; Kristensen et al. 2022). The simplest represent the ocean as a two-dimensional barotropic field with a free surface. This is also how we model the water level in this study. As Kelvin waves propagate as shallow-water waves whose phase and group speed are $\sqrt{gH}$, where $H$ is the local depth and $g\approx 9.81\,\mathrm{m\,s^{-2}}$ the earth’s gravitational acceleration, the amplitude and propagation speed of storm surges is sensitive to errors in both the gradient and the mean level of the bathymetry.
The availability of storm surge hindcasts for the Norwegian Sea, the North Sea and the Barents Sea region with sufficient spatial resolution, and spatiotemporal coverage is scarce, as shown by Fernández-Montblanc et al. (2020). Long hindcast integrations are necessary to assess water level extremes in coastal regions. To do this realistically requires high-quality and high-resolution atmospheric forcing and sufficiently high spatial resolution for the storm surge model. Storm surge hindcast archives serve mainly three practical applications. The first is to properly account for the overall statistical distribution of the water level, and the other is to assess the extreme value distribution, i.e., the return values for extreme high and low water level. In addition to these statistical measures, the scarcity of observational records means that accurate reconstructions of past events are of societal value in themselves. High-quality hindcast archives are also essential for establishing a baseline against which we can estimate the future projections of changes to the storm surge climate (Bernier et al., 2024). The objective of this article is to present such a open and freely available historical hindcast and to assess its validity for extreme value estimation of water level in the North Sea, the Norwegian Sea and the Barents Sea.
@@ 2 Observational dataset
Water level gauges from the coast of Norway (stations with NO prefix, also used in the study by Tedesco et al. 2024), Norwegian oil platforms (WL), Denmark (DK), Sweden (SW), Germany (DE), the Netherlands (NL), and the United Kingdom (UK) make up our observational data. Unfortunately, the license agreements do not allow open redistribution of the data we gathered. However, we share the code used for parsing and assembling the data from these sources into a single dataset (see Appendix A). We also summarize the source of the data from each country in Appendix B.
The data provided come from a variety of water level gauges. The location of the gauges and the associated measurement stations are shown in Figure 1. The time spans of the time series vary from station to station (see Figure 3b). Some stations have continuous recordings going back to 1915 (for some of the UK stations), though our dataset is built starting from 1979 onwards. As a consequence of the large time span covered, the kind of gauge, temporal resolution, and water level resolution of the measurements is not homogeneous in time even sometimes for a single individual station. Moreover, some data sources provide a quality flag as part of the data downloaded (depending on the data source and station, this may be either an automatic quality flag, or a human determined expert quality flag), while others only provide raw data without quality flag. We have, therefore, performed a detailed quality control cycle on all the data, which includes quality checking of individual measurements through both flag checking and statistical analysis, interpolation on a common time base, linear detrending, visual inspection, automated outlier detection, and a series of quality checks on running windows of observational data. The quality control removes approximately $9\%$ of the total observation data. The technical details of how this was performed are presented in Appendix B.
A tidal analysis was performed with the CNES pangeo-pytide python package  [^1] https://github.com/CNES/pangeo-pytide in order to estimate the tidal signal at each station and to detide the observed water level. For this, the time series obtained for each station is used to obtain the modal coefficients for the default 67 tidal modes used by pytide 222the corresponding modes are: O1, P1, K1, 2N2, Mu2, N2, Nu2, M2, L2, T2, S2, K2, M4, S1, Q1, Mm, Mf, Mtm, Msqm, Eps2, Lambda2, Eta2, 2Q1, Sigma1, Rho1, M11, M12, Chi1, Pi1, Phi1, Theta1, J1, OO1, M3, M6, MN4, MS4, N4, R2, R4, S4, MNS2, M13, MK4, SN4, SK4, 2MN6, 2MS6, 2MK6, MSN6, 2SM6, MSK6, MP1, 2SM2, Psi1, 2MS2, MKS2, 2MN2, MSN2, MO3, 2MK3, MK3, S6, M8, MSf, Ssa, Sa, and to perform tidal prediction on the same time base as the observations. By subtracting the harmonic tide analysis estimates obtained with pytide from the in-situ observations, we compute the (non-tidal) weather-induced residuals that are used to evaluate the storm surge model quality. This is further illustrated in Appendix B.
@@ 3.1 The ocean model
The Regional Ocean Model System (ROMS, see Shchepetkin and McWilliams 2005) is used for operational storm surge forecasting for Norwegian waters, as described by Kristensen et al. (2022). The model is set up to run in two-dimensional barotropic mode and the model domain is depicted in Figure 2. The governing model equations are thus the shallow water equations as described in Haidvogel et al. (2008). Let the total height be $h=H+\zeta$, where the $H$ is the equilibrium depth (Mean sea level is chosen as reference level in our simulations) and $\zeta$ is the sea surface deviation from the mean. By averaging the velocity over the total height $h$ we get (Kristensen et al., 2022; Tedesco et al., 2024),
@@ 4.1 Model performance evaluated against observations
The hindcast dataset has been extensively evaluated and validated against the observational dataset obtained in Section 2. In all the following analyses, the mean sea level for the entire hindcast period for each grid point has been subtracted from the modeled water level in order to establish mean sea level as the reference level. In order to establish a measure for the storm surge observation, the tidal predictions are subtracted from the time series for total water level for each of the stations. As mentioned in Section 3.1 and by Idier et al. (2019) and Horsburgh and Wilson (2007), the "pure weather effect" can not be completely separated from the total water level due to non-linear interactions between different factors contributing to the total water level at the coast. Hence, these non-linear effects are one of the possible sources of errors in all the following analyses. The work performed in the present study does not allow us to quantify the contribution to the total error from the different sources.
@@
Figure 3: Left panel: Comparison of standard deviation for storm surge between hindcast and observed time series for each station. In addition, the stations that were disregarded in the analysis are marked by triangles. Right panel: Difference between model and quality controlled observations scaled by the standard deviation for storm surge for each station for all stations for the entire hindcast period. The colorbar indicate differences in the range ±$\pm 1$ standard deviation. Note that the temporal coverage of the observing stations varies significantly. Stations are grouped by country.
@@
Figure 3a shows a comparison of the standard deviation for the time series for both the hindcast data and the observational data for all stations. This can be viewed as a measure of storm surge activity, and clearly shows that the hindcast data exhibit similar properties with regards to activity as the observations, and hence confirms the general quality of the hindcast. In Figure 3b we present the difference between the hindcast and observational data scaled by the standard deviation for storm surge for each station. This can be seen as a measure of the relative error, comparable between the different stations.
$\mbox{MAE}=\frac{\sum_{i=1}^{n}\mid\mbox{obs}_{i}-\mbox{mod}_{i}\mid}{n}$ (4)
$\mbox{RMSE}=\sqrt{\frac{\sum_{i=1}^{n}(\mbox{obs}_{i}-\mbox{mod}_{i})^{2}}{n}}$ (5)
$\mbox{MAE}/\sigma=\frac{\sum_{i=1}^{n}\mid\frac{\mbox{obs}_{i}-\mbox{mod}_{i}}{\sigma}\mid}{n}$ (6)
$\mbox{RMSE}/\sigma=\sqrt{\frac{\sum_{i=1}^{n}(\frac{\mbox{obs}_{i}-\mbox{mod}_{i}}{\sigma})^{2}}{n}}$ (7)
MAE [cm] RMSE [cm] MAE/$\sigma$ RMSE/$\sigma$
All data 9.7 12.4 0.42 0.54
Masked less than +/- 10cm 10.1 12.9 0.44 0.56
Masked less than +/- 30cm 11.8 15.2 0.52 0.67
@@
Table 1: General values for Mean Absolute Error (MAE), Root Mean Square Error (RMSE), MAE and RMSE scaled by the standard deviation ($\sigma$) of the observed storm surge for each station (MAE/$\sigma$ and RMSE/$\sigma$) averaged over all stations over the entire hindcast period. The rows with "masked" values indicate MAE and RMSE values when masking the data for the time periods when the observed absolute values of storm surge is less than $10$ and $30$ cm. The masking thresholds removes $43\%$ and $85\%$ of the data, respectively.
@@
Average values for Mean Absolute Error (MAE, see Equation 4) and Root Mean Square Error (RMSE, see Equation 5) for all stations combined given both in cm and as fractions scaled by the standard deviation of the observed storm surge for each station (MAE/$\sigma$, see Equation 6, and RMSE/$\sigma$, see Equation 7), averaged over the entire hindcast period, are presented in Table 1. When including all the available quality controlled observations in the comparison with the model data we get a MAE of $9.7$ cm and RMSE of $12.4$ cm, and a scaled error for MAE/$\sigma$ of $0.42$ and RMSE/$\sigma$ of $0.54$. The value for RMSE in the current work is in line with the work by Fernández-Montblanc et al. (2020) where they report an RMSE for the North Sea and Norwegian sea regions of $14$ cm and $11$ cm respectively. On average, the storm surge contribution to total water level is relatively small, as can be seen in the histograms in Figure 4, showing the distribution of modeled and observed water level, and in Figure 5, which displays the density scatterplot of observations vs modeled values. To investigate the events when there is actually a significant contribution from storm surge to the total water level, we have also added the values of MAE and RMSE where we have masked the data when the observed absolute value of the amplitude of storm surge is less than $10$ cm and $30$ cm in Table 1. These thresholds remove $43\%$ and $85\%$ of the data, respectively. As expected, and as explained by Kristensen et al. (2022), when the amplitudes of storm surge increase, the MAE and RMSE are expected to increase.
@@
The distributions of modeled and observed water level in Figure 4 indicate that the hindcast has too many occurrences of the values between $-50$ and 00 cm and from $20$ cm upwards, and too few in the range between 00 and $20$ cm when considering all stations combined (Figure 4a). For the Norwegian stations, the increased count for hindcast values is visible for the range $-25$ to 00 cm and $25$ to $75$ cm, and the most pronounced increase in count of observed values is found between 00 and $25$ cm (Figure 4b). For the range extending further above $+25$ cm, the hindcast has a slight increase in count for values up to about $+200$ cm for all stations. When examining the details of the tail of the distribution, for both positive and negative water level in Figures 4c and 4d, we note that the hindcast generally does not contain the same extremal values as the observations. Also, we would like to point out the four outliers with values below $-400$ cm and above $+400$ cm in Figure 4c. These are believed to be unphysical outliers in the observational dataset that evaded our quality control.
@@
Together with the previously mentioned non-linear interactions between e.g. tides and storm surge that are part of the observational data, but not accounted for in the model, there are other factors that can contribute to the total error described in this section. Other sources of errors include local contributions to the water level by wave setup (Stockdon et al., 2006; Melet et al., 2018), errors in the atmospheric forcing itself or how the momentum transfer from the atmosphere to the ocean is parameterized as stresses. The latter includes how momentum is transferred from the atmosphere to the ocean, and distributed in time and space, via the generation and decay of surface waves (Saetra et al., 2007; Breivik et al., 2015). Choices made during the model setup, like the parameterization of e.g. bottom drag, and selection of parameters can produce errors and inaccuracies. In addition, ice covered waters, where and when present, will experience a reduced momentum transfer from the atmosphere to the ocean (Wang and Bernier, 2023), which is also not accounted for in the current configuration of our hindcast. Lastly, we point out the model resolution as a possible source of error due to limited abilities to resolve features like e.g. coastline geometry and bottom topography.
@@
The heatmaps in Figure 5 display the two-dimensional distribution of corresponding hindcast and observation values. We note a small tendency for observations to be clustered around zero in the distribution for all stations (Figure 5a). The same is seen, although less pronounced, for the hindcast, with a slight increase in density along the $x$-axis. This could be caused by either unrealistic or erroneous observations still left in the dataset after quality control, or wrong characteristics for the hindcast for some stations/periods. The same is not seen in the distribution for the Norwegian stations (Figure 5b). Further investigation into the subject narrows the feature down to stations located in the Netherlands (most pronounced), Germany and Denmark. Since the number of data points with this behaviour is relatively small [^3] Further analysis show that the cases when the absolute value of the observation is below $10$ cm, and the absolute value of the model is larger than $75$ cm, account for 0.04% of the total quality controlled observations. (note the log scale for density), further efforts to identify and possibly remove these data points/stations has not been made. This might also be an indication that the NORA-Surge hindcast in some cases has problems capturing the exact behaviour of the storm surge for shallow stations located in the German Bight and surrounding area where tide-surge interaction might be of importance.
MAE [cm] RMSE [cm] MAE/$\sigma$ RMSE/$\sigma$
Denmark All data 10.7 13.7 0.33 0.42
Masked less than +/- 10cm 10.9 14.0 0.33 0.43
Masked less than +/- 30cm 12.2 16.0 0.38 0.49
Germany All data 13.9 17.8 0.39 0.49
Masked less than +/- 10cm 14.3 18.5 0.40 0.51
Masked less than +/- 30cm 16.2 21.1 0.45 0.59
Netherlands All data 11.7 15.0 0.44 0.57
Masked less than +/- 10cm 12.0 15.4 0.46 0.59
Masked less than +/- 30cm 13.8 17.8 0.53 0.68
Norway All data 6.9 8.6 0.44 0.54
Masked less than +/- 10cm 7.2 8.9 0.46 0.56
Masked less than +/- 30cm 8.5 10.4 0.53 0.65
Sweden All data 6.5 8.3 0.34 0.43
Masked less than +/- 10cm 6.8 8.7 0.35 0.45
Masked less than +/- 30cm 8.2 10.3 0.43 0.53
UK All data 8.5 10.7 0.46 0.59
Masked less than +/- 10cm 9.1 11.5 0.50 0.63
Masked less than +/- 30cm 11.2 14.0 0.64 0.80
Oil platforms All data 5.7 7.6 0.42 0.55
Masked less than +/- 10cm 6.4 8.7 0.46 0.63
Masked less than +/- 30cm 9.9 14.9 0.72 1.07
@@
As previously mentioned in Section 3.1, our hindcast model setup is based on the operational forecast model for storm surge used by the Norwegian Meteorological Institute (as described in Section 3.1). The model has been used for many years to provide forecasts of storm surge along the Norwegian Coast. However, even if the model domain covers a significant area outside of the Norwegian Coast, little attention has previously been given to the quality of the forecast for those areas. Geographical differences in the quality of the hindcast could mainly, but not limited to, be attributed to two factors: One is the fact that storm surges behave differently in different geographical areas due to, e.g., different characteristics of bottom topography, coastline geometry and weather patterns. The other could be that the model has worse performance and quality in some regions than others. The present work is not aimed at, or sufficient for, distinguishing between the two. Some geographical differences in the behaviour/characteristics of the storm surge can be seen in Figure 3a where we display the standard deviation for the model and observations for all stations (grouped by country). Figure 3b depicts the simple difference between model and observation over time as a fraction of the standard deviation for each station (as shown in Figure 3a). The figure show some differences between the individual stations, and perhaps some tendency towards systematic differences between the different countries, but these are not very clear. In Table 2 we provide the same type of comparison as in Table 1, but grouped by country. This emphasize that there are large variations in the quality of the hindcast when considering absolute values of MAE and RMSE, but when considering the "relative error" (MAE and RMSE based on differences scaled by the standard deviation) the geographical differences get less pronounced. The scaled values for MAE range from $0.33$ to $0.46$ and RMSE from $0.42$ to $0.59$ standard deviations. The smallest absolute errors are found at the stations in Sweden, Norway and the offshore oil platforms located in the North Sea, with values of MAE and RMSE, respectively, of less than $7$ cm and $9$ cm. This is less than the averages over all stations, and more in line with the values of MAE and RMSE of $3$ and $5$ cm for the errors reported at initial time ($+0$ h forecast lead time) by Kristensen et al. (2022) for the Norwegian storm surge forecast system evaluated at Norwegian stations. However, we would like to point out that the forecasts used in the evaluation by Kristensen et al. (2022) are post-processed and subject to correction by using the difference between model and observations, and hence exhibit lower error values than the free run in the NORA-Surge hindcast. The purpose of such corrections is to reduce low frequency error components in the reference level for the water level present in the operational forecast for a discrete set of stations on the Norwegian Coast. The correction does not affect the amplitude of storm surges, and does not correct the two-dimensional model field. Thus, such a correction is not applicable along the whole coastline in the present hindcast.
@@
We note two outliers in the Taylor diagram, the top left NO station with a correlation of approximately $0.7$ and the UK station with a correlation of $0.75$. The NO station is Ny-Ålesund (NO_NYA) located on the Arctic island of Spitsbergen, and the UK station is Newhaven (UK_Newhaven), located just west of Dover (in the eastern part of the English Channel). For the NO_NYA station, we note that there seems to be a negative trend in the difference between hindcast and observation in Figure 3b. This trend is explained by the rapid vertical land uplift at Svalbard (Kierulf et al., 2021). The UK_Newhaven station is placed near the south-western boundary of the hindcast model domain, and the fact that it is located west of Dover means it will be subject to storm surges that enter the English Channel from the west. The lack of proper boundary values from an outer model, as mentioned in Section 3.1, could cause some of the storm surge signal that should have entered from the North Atlantic to be missing in the hindcast for this area of the English Channel.